OpenStudy (bookworm14):

Could someone please explain how to work this out step by step? It is easy, but I am getting confused. THANKS! ~Bree x^2 - 10xy + 24y^2

OpenStudy (anonymous):

so what is the question, baby?

OpenStudy (bookworm14):

it is posted above, but here it is :\[x^2 -10xy+24y^2\]

OpenStudy (anonymous):

what do you have to do with that function?

OpenStudy (anonymous):

graph? factor? solve for x?

OpenStudy (bookworm14):

I have to factor it

OpenStudy (anonymous):

ok

OpenStudy (bookworm14):

I forgot how to do it with two variables lol

OpenStudy (anonymous):

Simplifying x2 + -10xy + 24y2 Reorder the terms: -10xy + x2 + 24y2 Factor a trinomial. (x + -6y)(x + -4y) Final result: (x + -6y)(x + -4y)

OpenStudy (bookworm14):

how do you factor it though? im not used ot have multiple letters...

OpenStudy (anonymous):

let me work it out! one sec

myininaya (myininaya):

do you know how to factor u^2-10u+24?

OpenStudy (bookworm14):

yes

myininaya (myininaya):

what do you get when you factor it

OpenStudy (bookworm14):

(u-4)(u-6) ?

myininaya (myininaya):

x^2-10xy+24y^2 is the same way you will have (x-4y)(x-6y)

myininaya (myininaya):

x^2+Bxy+Cy^2 if factorable will look something like this: (x+ny)(x+my)

OpenStudy (bookworm14):

ok, i think I see what you are talking about. so i just have to insert the "y" into both sets of parenthesis?

myininaya (myininaya):

yeah next to that last number in each so when you multiply the last times the last you will get -4y(-6y)=24y^2 which is the last term in your expression

OpenStudy (bookworm14):

ohhhh, thanks! I am pretty sure i understand :)

myininaya (myininaya):

and when you do first times the last from both that is you do x(-6y)+(-4y)x=-6xy-4xy you get -10xy :)

OpenStudy (bookworm14):

i get it! thank you so much!! :D

myininaya (myininaya):

which is your middle term in this case

OpenStudy (bookworm14):

the middle term would be -10xy right? i might have this wrong :/

myininaya (myininaya):

yes the middle term is -10xy :) other examples: \[x^2+5xy+6y^2 =(x+2y)(x+3y)\] \[x^2-4xy+3y^2=(x-3y)(x-1y) \text{ or } (x-3y)(x-y)\]

OpenStudy (bookworm14):

thank you :) im taking a screen shot for future reference lol :)